If the median AD of triangle ABC be perpendicular to AB then prove that tan A + 2tan B = 0
(I think in above problem it should be BC in place of AB)


Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Yes it should be Bc rather than AB..

Once you do this correction, then it becomes a simple question...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

gautam gangadar
6 years ago
Ans is zero it can be proved using sine rule in the two triangles triangle ADC anf Triangle ABCin tri ABC,. (SinA )/BC. = (SinB ) / ACin tri ADC , (Sin (A-90))/DC = (Sin(B+90))/ACsolve it you get the answer.